The manufacturer of a brand of juice drink claims that the mean content of its product is 210 ml with a standard deviation of 6.7 ml.



What is the probability that a randomly selected sample of the product contains more than or equal to 213 ml?



What is the probability that a sample contains less than 200ml?

1
Expert's answer
2021-12-20T03:59:40-0500

Let "X=" the volume of juice in a sample: "X\sim N(\mu, \sigma^2)."

Given "\mu=210\ ml, \sigma=6.7\ ml."

i)


"P(X\geq 213)=1-P(X<213)"

"=1-P(Z<\dfrac{213-210}{6.7})"

"\approx1-P(Z<0.44776)\approx0.32716"

ii)


"P(X<200)=P(Z<\dfrac{200-210}{6.7})"

"\approx P(Z<-1.49254)\approx0.06778"